A and B are the point (3,-1) and (2,3) respectively. l is the line 8y - 2x - 3 = 0 show that l passes through the midpoint of AB.
Answer by mananth(16946)

(3,-1) and (2,3)
The mid point will be (3+2)/2 = (5/2)
(-1+3)/2 = 2/2 = 1
The co-ordinates of mid point are ( (5/2),1 )
If this lies on the line 8y - 2x - 3=0
It must satisfy the equation
Plug the co ordinates of the mid point
8*1 -2*(5/2)-3 =0
LHS = RHS
so the midpoint lies on the line